OpenStudy (anonymous):

Eighty ounces of Lazy Lawn fertilizer covers 1,250 square feet of lawn. a. How many ounces would be required to cover a 4,000-square-foot lawn? Do not enter units in your answer. ounces b. If Lazy Lawn costs $1.19 for a 32-ounce bag, what is the total cost to fertilize the lawn? $

OpenStudy (anonymous):

1250/80 = 4000/x solve for x number of bags = x / 32 cost = 1.19 times number of bags

OpenStudy (anonymous):

Okay, this is for a. You can find how much it would require by figuring out how many times 1,250 goes into 4,000. Multiply the number you get by 80 (the amount of fertilizer to cover 1,250) to find out how much fertilizer you need to cover 4,000 square feet. I don't explain well XD